OpenStudy (anonymous):

express the ratio 45 seconds to 30 min as fraction in simplest form

OpenStudy (anonymous):

Ok in this question we will jhave too convert wither of the quantity into others specs for here I will convert minutes into seconds(just to make it length and easy): \[\frac{ 45 \sec }{ 30 \min }= \frac{ 45 \sec }{30 * 60 } = \frac{ 3 }{ 2 * 60} =\frac{ 1 }{ 2 * 20 }= \frac{ 1 }{ 40}=1:40\]

OpenStudy (anonymous):

@briannafeet your question's answered

OpenStudy (mokeira):

first change min to sec \[30 \times 60= 1800mins\] now you can continue like this \[\frac{ 45 }{ 1800 }\] then simplify the fraction to get \[\frac{ 1 }{ 40 }\] = 1:40

OpenStudy (mokeira):

1800 sec...not 1800mins..typing error

Elsa213 (elsa213):

you can't take ratio of 2 quantities with different units :) here, we NEED to convert both time into seconds and only then take ratio. 45/1800=....

Elsa213 (elsa213):

then the answer is when you simplify it 5 over 200 then again is 1 over 40

Elsa213 (elsa213):

i got this from: http://openstudy.com/study#/updates/533ebcd7e4b02f8a3bd4a48a